OpenStudy (anonymous):

Juanita and Keenan own a camping supply store and just put in an order for flashlights and sleeping bags. The number of flashlights ordered was three times the number of sleeping bags. The flashlights cost $12 each and the sleeping bags cost $45 each. If the total cost for flashlights and sleeping bags was $1215, how many flashlights and how many sleeping bags were ordered. So I know I need to have a system of equations here. My first equation is 12x + 45y = 1215. However, I can't figure out the second. Any help would be much appreciated!

OpenStudy (anonymous):

flashlight = x sleeping bag = y 1215 = 3(12)y + 45y

OpenStudy (anonymous):

x = 3 times the number of sleeping bags

OpenStudy (anonymous):

But don't you need a system of equations to be able to solve for x and y?

OpenStudy (anonymous):

not sure. 1215 = 36y + 45y 1215 = y(35 + 45) 1215 = 80y 1215 = y

OpenStudy (anonymous):

sorry 1215/80 = y

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

thats not a whole number

OpenStudy (anonymous):

because i added them wrong? lol

OpenStudy (anonymous):

I think if we determine that x = flashlights and y = sleeping bags then we need to solve for each variable. I was thinking 12x + 45y = 1215 3x = y (because you purchase three times more flashlights than you did sleeping bags.

OpenStudy (anonymous):

y = 15, 3 times x = 3y = 45. 15(45) + 45(12) = 1215

OpenStudy (anonymous):

for simultaneous equations you would need two cases. say "yesterday he sold twice as many as today" aswell. then you could set up one equation for yesterday, and one for today

OpenStudy (anonymous):

but because you have one variable relating to another "3times more" in this one case. its going to be an equation where you replace a variable.

OpenStudy (anonymous):

Yes, I think I have it because when I plug x and y back in the add up to 1215.

OpenStudy (anonymous):

yep :)